### v2

# Implementing Fairness in AI Classification: The Role of Explainability

Abstract
In this paper, we propose a philosophical and experimental investigation of the problem of fairness in AI classification. We argue that implementing fairness in AI classification involves more work than just operationalizing a fairness metric. It requires establishing the explainability of the classification model chosen and of the principles behind it. Specifically, it involves making the training processes transparent, determining what outcomes the fairness criteria actually produce, and assessing their trade-offs by comparison with closely related models that would lead to a different outcome. To exemplify this methodology, we trained a model and developed a tool for disparity detection and fairness interventions, the package FairDream. While FairDream is set to enforce Demographic Parity, experiments reveal that it fulfills the constraint of Equalized Odds. The algorithm is thus more conservative than the user might expect. To justify this outcome, we first clarify the relation between Demographic Parity and Equalized Odds as fairness criteria. We then explain FairDream’s reweighting method and justify the trade-offs reached by FairDream by a benchmark comparison with closely related GridSearch models. We draw conclusions regarding the way in which these explanatory steps can make an AI model trustworthy.
Keywords: Explainable AI ; Fairness ; Equalized Odds ; Demographic Parity; Calibration; GridSearch ; Trust
1 Introduction
Determining whether an algorithm provides ‘fair’ predictions is a challenging task, although one that is of central practical and ethical importance for society. The difficulties pertain in part to the fact that different notions of fairness exist and compete with each other. On the theoretical side, plethora of statistical metrics exist to quantify the extent to which the predictions of an AI model meet the user’s expectation on fairness, in particular regarding a minority that must not be harmed, [Hellman, 2020] [Mehrabi et al., 2021] [Wan et al., 2023]. Yet, attempts to compare these algorithmic fairness metrics face impossibility theorems, implying that the various fairness criteria proposed cannot be jointly satisfied in all cases (viz. [Kleinberg et al., 2016, Hellman, 2020, Hedden, 2021]).
As an illustration of this, some concrete cases have shown us that the recommendations of an AI may be viewed as fair by its designers, but as unfair by external evaluators. The case of COMPAS (Correctional Offender Management Profiling for Alternative Sanctions), an algorithmic tool intended to score recidivism risk, is a telling example. Notoriously, the ProPublica 2016 report accused COMPAS of racial bias [Larson et al., 2016], by showing a higher rate of false positives in African Americans for “high-risk” scores and a higher rate of false negatives in Caucasians for “low-risk” scores. In response, the designers of COMPAS (Northpointe) replied that ProPublica artificially built these categories (by setting a boundary at 5/10 of recidivism risk), and that for each different risk level, the statistical performance of COMPAS (as measured by ROC-AUC, i.e. rate of true positives compared to false positives) was actually the same across both groups, and even slightly more favorable to black defendants [Dieterich et al., 2016].
But plurality of metrics is arguably not the main challenge for fairness. A more common obstacle is opacity concerning the mechanism behind a classificatory verdict. This concerns cases in which the decision taken by an AI or by a human is issued without the user having any idea of the principles behind it. Examples could be multiplied: why did the system not give me the option to go to Uni when I thought I had the credentials? (see [Grosperrin, 2023]). Why was my loan refused when I think I could refund it? ([Purificato et al., 2023]) Although no classification scheme may be immune to error, we believe that errors can at least be mitigated when the user has access to the principles behind the decision, if only because they can question them and suggest amendments.
In this paper, we propose to look at the problem of defining a transparent and explainable AI system for inequality detection and correction. As with other approaches to AI fairness developed in diverse domains, including medicine [Chen et al., 2023], law [Lagioia et al., 2023], or mortgage lending [Lee and Floridi, 2021], we anchor our philosophical analysis to a particular case study, to draw general lessons from the way in which bias can be detected and corrected using AI. Specifically, we trained an AI model and developed our own fairness package (FairDream) to detect inequalities and then to correct for them. As experiments with FairDream reveal, however, the settings proposed by the system don’t necessarily produce expected outcomes, and so they call for an explanation in order to make the system not just usable, but also transparent, and trustworthy.
The way FairDream works is by implementing certain criteria for fairness correction. In relation to that, we draw attention to a tension between two fairness criteria: one promoting the same overall positive rate across groups (Demographic Parity), another recommending to equalize true positive rates and false positive rates across groups (Equalized Odds). In the literature on algorithmic fairness, Equalized Odds is also opposed to the notion of Calibration ([Mayson, 2018, Long, 2021, Hedden, 2021, Barocas et al., 2023]), often leaving Demographic Parity as a distant contender. However, our reasons to focus on the opposition between Equalized Odds and Demographic Parity rather than with Calibration are twofold.
First of all, demands of Demographic Parity are real in various segments of society, in relation to gender, age, or ethnicity. Demographic Parity has thus been interpreted as a legal requirement, e.g. on EU Non-Discrimination Law [Wachter et al., 2020] or for the risk assessment mandated by Pennsylvania [Selbst et al., 2019]. Secondly, while we trained an AI model to enforce Demographic Parity, it eventually converged on the enforcement of Equalized Odds instead, in ways that are not just coincidental. Our experiments reveal that it is a property of FairDream to fulfill fairness objectives which are conditional on the ground truth (Equalized Odds), even when the user wants to achieve unconditional equality (Demographic Parity). While this may be seen as an anomaly, we first explain this property of FairDream, and then we use it to propose an argument for fairness metrics conditioned on true labels. FairDream is intended to enable a user to perform fairness interventions. We explain the way in which the criterion of Equalized Odds constrains such interventions, in particular by comparing it to similar in-processing methods that forego this constraint. From this comparison, we draw the lesson that one way of making an AI system trustworthy is to provide a comparison and contrastive explanation of its results with minimally modified versions of the system that produce significantly different outcomes.
In Sections 2 and 3, we first present our training data and the package FairDream. Section 4 introduces the specific property of FairDream of enforcing Equalized Odds, and Section 5 proposes to explain this property by comparison with a closely related approach, the so-called GridSearch method, which can diverge from true labels in order to enforce Demographic Parity. In Section 6, finally, we present some arguments in favor of a metric conditioned on true labels. Firstly, we discuss how FairDream performs in terms of calibration, and we explain calibration is not a relevant criterion from our perspective, including when it is interpreted as a requirement of equal precision relative to a threshold. Then, we respond to the suspicion that the criterion used might be too conservative. We argue that a more revisionary notion of fairness like Demographic Parity too can be enforced, but provided its effects are carefully controlled for. Finally, in Section 7 we conclude with broader lessons regarding the way in which the explanatory steps taken here make the model trustworthy.
2 Age disparities in the evaluation of income
Predicting income matters in situations in which investors need to make funding decisions and ensure their customer’s capabilities. Depending on how they situate the customer’s assets, such predictions may be used in banking as a basis to grant credits, in wealth management to recommend investment products, but also in recruitment to set salaries and make attractive offers [Meng et al., 2018].
For our experiment, we thus manipulated the well-known Census dataset, which aggregates data from 48,842 US citizens in 1994.111https://www.kaggle.com/datasets/uciml/adult-census-income. The dataset includes 14 characteristics such as age, work class, years of education, marital status, and geographical origins. A fifteenth column, finally, displays a binary encoding of whether the agent earns less or more than \$50,000 a year (Cf. Figure 1).
Figure 1: A sample of the Census dataset
The task for an algorithm is to draw inferences from the dataset, namely to classify whether agents earn over \$50,000 with the highest possible accuracy. [^2] Throughout this paper, we use accuracy in the sense of predictive performance, as measured by the ROC-AUC. Based on the patterns of Census, the model has to capture the correlations between occupation, age, or education level, and the target (whether the income is above \$50,000). For such a statistical learner, the goal is to be able to generalize to new data which encompass only the fields of age, education level, etc.
In the case at hand, the distribution of ages appears as a good predictor for the income variable (Cf. Figure 2). The proportion of individuals earning over \$50k in the group under 29 years of age (brown area/blue area) is very small (453/11,295 individuals, i.e. 4%). However, this proportion grows quite significantly between 29 and 37 years of age (2393/9914 individuals, i.e. 24%). Therefore, a learner might use age (among other features) as a critical factor in order to predict the wealth of individuals – in particular, to generalize its estimates of income to data with unknown labels.
Using the Census dataset, we thus built a machine-learning model whose purpose is to minimize the error between its predictions and the actual income of people. Because the practitioner can choose different learners and different metrics of statistical performance, we built a tree-based AI model with the XGBoost library and evaluated it using the standard ROC-AUC measure. The models of trees ensembles as their evaluation through ROC-AUC are, indeed, broadly used in binary classification on tabular data [Shwartz-Ziv and Armon, 2022]. [^3] An XGBoost classifier, – for eXtreme Gradient Boosted trees,– is an ensemble of trees. Each tree tries to find the best features (splitting populations on their values) to distinguish between individuals earning more or less than \$50,000. Boosting relies on a sequence of weak trees, progressively correcting each others to reach the best predictor. In XGBoost, the computation is made faster through gradient approximation of the errors of the trees, and techniques such as tree-pruning are used to find new trees faster [Chen and Guestrin, 2016]. For stability of the following experiments, we reduced the architecture of the gradient boosted trees to a minimal tuning. By default, we set the model to 1,000 estimators, the maximal depth of each tree being 3 splits on features.
Based on the distribution of income by age, we can expect our tree-based model to rely on age to classify clients in terms of predicted income. A problem can appear, however, if the user judges that age differences are not a fair basis for credit allocation, product recommendation, or wage offer. This could happen due to regulatory requirements [Adams-Prassl, 2022], or if a bank or a wealth manager is committed to policies favorable to youth [Calisir and Gumussoy, 2008]. In those cases, the tendency of a classifier to align with data distribution, and to use age in order to discriminate among clients, might be seen as a form of unfair discrimination. The challenge of our case-based approach is therefore to explore this tension between the model’s tendency to discriminate on the basis of a salient correlation, and the fairness requirements of a business user.
Figure 2: Proportion of individuals earning over \$50,000 by age. The red and green lines delineate the groups 17-29 and 29-37 between which FairDream detects a disparity.
3 The package FairDream
How can we assess whether an AI model in charge of predicting level of income is fair before deploying it? To answer this question, we implemented our own algorithmic fairness package (FairDream). In FairDream, the main purpose is to give lay users, without any background in data-science or in theories of fairness, an insight into the baseline AI model, and subsequently, a handle on the process of correction.
To enable the users’ understanding, we started the correction with the simplest metric of FairDream, the overall percentage of individuals predicted as positive. This metric basically aims for Demographic Parity [Dwork et al., 2012], which means that it offers to equalize the percentage of individuals selected by the model as earning at least \$50,000 across groups, based on some given feature. [^4] As the detection of disadvantages is based on intervals of age, education level, etc., we adopted a group-based correction method. In this approach, either the law or the sensitivity of users defines one or several features (e.g. sex) on which groups or subgroups ought to be treated equally. There also exist individual-oriented approaches, which involve treating similar individuals in the same manner (according to a similarity distance, see [John et al., 2020]).
FairDream takes as input an already trained predictor (the baseline XGBoost we trained) and an indication of the users’ preference for fairness regarding the classificatory variable of interest (i.e. how much the user wants to impose parity regarding the percentage of individuals labeled as “\$$\geq\$50k$” by the XGBoost across groups). The procedure implemented in FairDream involves two steps, a step of detection and a step of correction.
Detection - The “Discrimination Alerts” algorithm detects disparities in how groups are treated. Each feature is inspected. If age intervals, certain jobs, or nationalities are under-selected by the model, a discrimination alert is issued. The algorithm hence makes users aware of disadvantaged groups, potentially distinct from the user’s beliefs on what those groups might be.
On the sample of Census used for training, FairDream  warns for a ratio of 1 to 5 in the model’s predictions: for example, for the 17-29 year-old individuals, the portion of the sub-population selected is 12%, against 66% for 29-37 year-old (Cf. Table 1), thereby triggering an alert.
Correction – With these alerts, the user can form a justified opinion about imbalance between populations; and they can decide which gaps between groups of features to correct for. In our experiment, we simulated the approach of a decision-maker whose normative preference is to reduce the difference between older and younger clients.
We thus passed the feature ‘age’ to FairDream, so as to build a model bridging the gap between ages. Five new XGBoost models were trained and put in competition to reach a similar statistical performance with the initial XGBoost, [^5] See Section 5.2 for more details on training FairDream models. while better classifying previously disadvantaged younger clients. Comparing the five new models with the baseline, FairDream selects the model satisfying the best trade-off between accuracy and fairness (see Figure 3). [^6] It is generally stated in the algorithmic fairness literature that an algorithm aiming to satisfy both fairness and statistical objectives can nothing but decrease in statistical performance [Kleinberg et al., 2016], compared with a perfectly-tuned model only paying attention to a statistical indicator (ROC-AUC). One challenge for the business-user is therefore to find an optimal trade-off for their use-case [Lee and Floridi, 2021].
In our experiment, the stat_score is the indicator of predictive performance chosen by the user, such as ROC-AUC. The fair_score is maximized when the indicators of fairness determined by the user are equalized across groups, such as the overall positive rates for Demographic Parity. We implemented it as a weighted linear sum of the differences of fairness indicators between groups. The best model is then selected as the one with the highest ${\tt trade\_off\_score}=1/3*{\tt stat\_score}+2/3*{\tt fair\_score}$, if for instance the user prefers fairness over accuracy.
Figure 3: Selection of the best corrected model (green), based on a combined measure of statistical performance and fairness.
Table 1 shows the predictions of the baseline model at the first step of the procedure, while Table 2 shows the predictions of the model selected as best by FairDream at the second step. Each table highlights the rates of positive predictions (overall positive, true positive, false positive) of the corresponding model’s predictions for two ages groups. The model is evaluated on each age group, first at the global level, then looking at the individuals having the true label ‘\$$\geq\$50,000$’, finally looking at the individuals having the true label ‘\$$\geq\$50,000$’.
At the global level, comparing Table 2 and Table 1 shows that the disparity between age groups in the corrected model is indeed less than 1 to 5, but the correction seems surprisingly small (16 % vs 59 % instead of 12 % vs 66 %), leaving a gap whose ratio is still of 1 to 3.7. When comparing by true labels, however, we do see an important correction, since true positive rates and false positive rates come out nearly equal (88% to 89%, 3% to 4%), unlike in the baseline model (68% to 90%, and 2% to 20%). This suggests that the model is not fulfilling Demographic Parity, but a different criterion, known as Equalized Odds. We review the difference between them in the next section.
Table 1: Overall Positive, True Positive, and False Positive rates in the baseline model
Table 2: Overall Positive, True Positive, and False Positive rates in the FairDream model
4 Demographic Parity vs Equalized Odds
Adopting the taxonomy of [Wachter et al., 2020], we may first interpret the desire to diminish the gap between older and younger clients as an unconditional wish, namely to predict equal proportions of 17-29 and 29-37 year-old as earning over \$50,000, regardless of their true level of income. That would be translated into an objective of “Demographic Parity” [Dwork et al., 2012]: [^7] We notate $Y$ whether an individual actually earns over \$50k, and $\hat{Y}$ the prediction of the model; $A$ and $B$ respectively represent whether an individual is in the 17-29-year old or in the 29-37 year-old category.
$p(\hat{Y}=1|A=1)=p(\hat{Y}=1|B=1)$
This objective of Demographic Parity is far from being achieved in the new model: only 16% of younger clients are selected against 59% for their elders. Although the ratio is a bit more advantageous (3.7 to 1 versus 5.5 to 1), the corrected model makes the same overall difference between groups of age as the baseline model did (Cf. Tables 1 and 2, red highlights).
Yet, the way age groups are treated by the FairDream  model leaves room for a different interpretation, based on their ground-truth label. The percentages of people predicted by the model as earning over \$50,000 can be compared between subgroups whose income indeed exceeds \$50,000 (true positive rates), or whose income is below \$50,000 (false positive rates). Determining if the model does better to classify them according to their true income corresponds to a fairness metric of “Equalized Odds”, namely equalizing true positive and false positive rates [Hardt et al., 2016, Barocas et al., 2023]:
$\left\{\begin{array}[]{ll}p(\hat{Y}=1|A=1\wedge Y=1)=p(\hat{Y}=1|B=1\wedge Y=1)\\ p(\hat{Y}=1|A=1\wedge Y=0)=p(\hat{Y}=1|B=1\wedge Y=0)\end{array}\right.$
Adopting this lens shows that younger clients are clearly reevaluated. We observe that the probabilities of being classified as earning at least \$50k are almost the same for those truly at that income level (88% vs. 89%) and for those who are not (3% vs 4%), in the younger as in the older age category (Table 2, green part). Whereas in the baseline model, there is a gap of almost 20% between these rates, disadvantaging younger clients (Table 1).
This is unexpected, since in this experiment we actually asked for the overall percentages of clients earning over \$50,000 to be predicted as equal across ages by the FairDream model. Mathematically, this objective corresponds to Demographic Parity. Yet, we see that Demographic Parity is not achieved by the FairDream  model. Instead, it actually produces a different result, namely Equalized Odds. Is it an accident? Or is it rather an emergent feature of the model, which FairDream can help us to rationalize? We proceed to clarify this issue in the next section.
5 Explaining FairDream’s results
To better situate the correction method of FairDream, we need to say more about the landscape of “algorithmic fairness”, that is about extant methods to bring a model closer to statistical fairness criteria [Mehrabi et al., 2021]. We describe the in-processing character of the method, and we explain the distinctive way in which reweighting works in it, compared to GridSearch methods that obey similar principles.
5.1 An in-processing method
Once we set the percentage of clients predicted as earning over \$50k to be equal across ages (Demographic Parity), the fairness techniques can equalize them before, during, or after training of the model [Alves et al., 2021]. Although they are implemented in diverse ways, these techniques mainly rely on the following architectural principles:
Pre-processing: Pre-processing methods intervene on the training data, to feed the future model with rebalanced or more favorable data for harmed individuals. The methods can change the labels (e.g. FairBatch [Roh et al., 2020]) or the frequency of minorities in the training dataset (e.g. Reweighing [Calders et al., 2009]).
Post-processing: Post-processing occurs after the model is trained. In classification, it takes place when we convert the probabilities into the predicted event. Some examples include adjusting the thresholds, which enables harmed individuals with lowest scores to be labeled as earning over \$50k ([Hardt et al., 2016]).
In-processing: In-processing methods generally consist in minimizing the loss function, subject to a fairness constraint, to satisfy a trade-off between the performances of the statistical and fairness metrics of the user [Wan et al., 2023].
FairDream is an in-processing method, in which the correction step takes place during model training. Underlying this choice, we considered that intervening outside the model unnecessarily increases the opacity of an AI system. In particular, the other two methods can lead the model to learn from unrealistic samples (in the case of pre-processing), or to generate new thresholds for the same group each time the model is deployed in a new context (case of post-processing).
5.2 Reweighting for transparent processing
In our attempt to make the method as transparent as possible to lay users, our method uses the principle of reweighting ([Calders et al., 2009]. Reweighting of harmed individuals generally matches the intuition of lay users concerning compensatory justice [Velasquez et al., 1990]. To wit, if the weight of an error made on a 28-year old is three times the weight of an error on older clients, the model has to fit its parameters while “paying three times more attention” at previously harmed individuals.
To make the correction of the model fit the lay intuition, FairDream  reweights groups in an ascending way. The weight of an error on a group $S_{k}$ grows with its previous disadvantage. The new sample weight $w_{n}(S_{k})$ of FairDream model #$\#n$ increases with the number of harmed individuals (linearly), and with the difference of fair scores in the baseline model (exponentially). [^8] See Appendix A for more details on FairDream’s correction step. For example, in FairDream’s model #$\#2$, which was selected as the best one, the weight of an error made by the classifier during training on 17-29-year old ($w_{2}(S_{1})$=0.8) became eight times the error on 29-37-year old ($w_{2}(S_{2})$=0.1).
[…]
$L(\phi_{\theta,n},\ W_{n})=\ \sum_{1\leq k\leq p}\ \sum_{i\in S_{k}}{w_{n}(S_{k})\ast\ l(y_{i},\widehat{y_{i}})+\Omega(\theta)}$ (FairDream Loss)
In FairDream, the correction for fairness is group-wise. Instead of “levelling down” the well off group to make its OPR closer to the worse off, say on sex, FairDream focuses on enhancing the OPR of women with equal or slightly lower OPR of men as in [Mittelstadt et al., 2023].
In our case, however, the selection of optimal classifiers is granted through group-wise reweighting during processing. [^10] To “level up” the worst treated group, [Mittelstadt et al., 2023] visualize by group the trade-off between the chosen selection rate and accuracy on a Pareto frontier, helping adjust thresholds for post-processing. As wished by [Castro, 2019], FairDream concretely implements costs of errors (FNR and FPR) proportionally to the structural disadvantage experienced by a group in society (e.g. women who need to be reevaluated on credit allocating, taking into account the disadvantages they face up in education, recruiting, income, etc.).
Hence, FairDream combines the optimality of “levelling up” with the transparency of the process of enhancing disadvantaged groups, through differentiated attention.
5.3 Experimental comparison with GridSearch
To investigate the proper effects of FairDream’s correction, we compared it in a benchmark experiment with a closely related fairness method: GridSearch [Agarwal et al., 2018], also an in-processing method of cost-sensitive classification. Described below and synthesized in the tables of Appendix B, the full results of our experiment are accessible in the GitHub benchmark repository: https://anonymous.4open.science/r/weights_distortion_impact-15/.
In GridSearch, the statistical performance and fairness tasks are split. More precisely, the loss is optimized under the fairness constraint $\mathcal{F}$, e.g. penalizing the model when the overall true positive rates are not equalized (with $\eta>0$):
$\min_{\theta}L(\phi_{\theta,n})\ \text{such that}\ \mathcal{F}(\phi_{\theta,n})\leq\eta$
The classifier $\phi_{\theta,n}$ has to find the parameters $\theta$ which minimize the accuracy loss $L(\phi_{\theta,n})$, so long as the fairness constraint $\mathcal{F}$ is under the threshold $\eta$. Which amounts to the loss function:
$L\left(\phi_{\theta,n},\ \lambda\right)=\ \sum_{i=1}^{m}\ l\left(y_{i},\widehat{y_{i}}\right)+\Omega(\theta)+\lambda^{T}\left(\mathcal{F}\left(\phi_{\theta,n}\right)-\eta\right)$ (GridSearch Loss)
Whereas FairDream  directly integrates the fairness objective through the sample weights $W_{n}$, GridSearch uses the Lagrange multipliers $\lambda$ to stress the violation of the fairness constraint. [^11] For more details on the GridSearch algorithm, see [Agarwal et al., 2018]. GridSearch starts with a grid of values $\lambda$ as FairDream  with the weights $W_{n}$, which convey in an accessible way the idea of a classification sensitive to protected individuals or not.
We conducted the experiment over multiple types of models. To grant stability of the experiments, we selected the event predicted by the model (earning over \$50,000 or not) for the threshold maximizing the F1-score, commonly used in machine-learning for imbalanced classification as in Census. [^12] The F1-score computes the harmonic mean between Precision and Recall for a given threshold. Hence, maximizing the F1-score helps satisfy the Precision / Recall trade-off. We used the Scikit-learn metric of F1-score.
  • Gradient boosted trees (using the XGBoost library: 1000 estimators, with the maximal depth of each tree being 3 splits on features) – to keep on investigating the initial type of model we used in Section 3.
  • Random forest trees (using the Scikit-learn library: 100 estimators, with the maximal depth of each tree being 3 splits on features) – to introduce a lighter tree-based model.
  • Neural networks (using the PyTorch library to build a sequential model alternating linear layers (14, 1000) $\to$ (1000, 230) $\to$ (230, 2) and ReLU layers to break linearity).
  • Logistic regression (using the Scikit-learn library, with the “liblinear” solver, performing approximate minimization along coordinate directions) – to introduce a simpler model in the benchmark.
For each type of model, a baseline model was trained with these default parameters, regardless of fairness objectives. Then to investigate the convergence of FairDream towards Equalized Odds, we analyzed the model through the lens of Demographic Parity, as if Demographic Parity was the initial fairness purpose set by lawmakers. When inequalities of overall selection were more than 3:1 across groups (e.g. eligible men for a loan = 42% vs 11% for women), we started a correction to mitigate gaps created by the model on that feature (e.g. sex).
With the same default parameters as the baseline model, GridSearch and FairDream tested new weights on 10 new models. The goal of the competing models was to simultaneously maximize a global statistical criterion (ROC-AUC) and equalize the overall positive rates across groups (Demographic Parity).
Our hypothesis was that if we set FairDream to equalize the groups across a fairness objective which is not conditional on true labels (in this experiment Demographic Parity), FairDream would nevertheless equalize the groups according to a fairness measure conditioned on true labels (here Equalized Odds).
After the best model of FairDream and the best model of GridSearch were selected, we found confirming evidence for our hypothesis (Cf. Appendix B). Let us illustrate this with one example from our experiments. GridSearch and FairDream tried to mitigate the gaps in overall positive rates between men and women generated by a baseline random forest (purpose of Demographic Parity). When we observe the new gaps in overall positive rates between the sexes, GridSearch has enhanced the position of women (now 100% predicted as earning over \$ 50,000). This is way more than in FairDream, which only selects 20% of the women (Cf. Figure 4).
However, the difference in true positive rates between GridSearch and FairDream is now a slight one (Cf. Figure 5, left). In FairDream, the percentage of individuals correctly predicted to earn over \$50,000 is now closer across the sexes (83% vs 76%, now slightly promoting women, where this difference is of 18% vs 80% in the baseline model). Likewise, false positive rates between male and female individuals are made closer than in the baseline model by FairDream(Cf. Figure 5, right: 13% vs 22%, where this difference is of 0.2% vs 24% in the baseline model). These results show that, even though we set FairDream to respect Demographic Parity, instead it achieved Equalized Odds by equalizing the true and false positives across populations, relative to what the baseline classifier does.
Obviously, the GridSearch model selects more women as true positives than FairDream does. However, this is at the expense of a drastic loss in statistical performance. To achieve equal positives rates, GridSearch actually predicted every woman to earn over \$50,000, making it like a random classifier (ROC-AUC = 56%, false positive rate = 100%).
On the contrary, FairDream bridges the gap between the sexes by improving on the accuracy of the baseline classifier (ROC-AUC = 80%). While the overall percentage of positives is not equally high in both sexes, it seems to achieve the maximal positive rate possible, according to the true labels. [^13] We explore this point later on, in Section 6.
The correction on the feature “sex” with random forest models to equalize the overall positive rates confirmed the tendency we observed across the results of our benchmark. In general, GridSearch better fulfills the fairness objectives, but at the expense of accuracy. On the contrary, keeping an accuracy comparable to the baseline model, FairDream generally performs better to equalize the true and false positive rates - even when it is not the fairness objective which was set to be maximized (see Appendix B and the GitHub benchmark repository for detailed visualizations: https://anonymous.4open.science/r/weights_distortion_impact-15).
Figure 4: Evaluation on Demographic Parity - Overall Positive Rates with FairDream, GridSearch and Baseline Random Forest Models on Sex
Figure 5: Evaluation on Equalized Odds - True and False Positive Rates with FairDream, GridSearch, and Baseline Random Forest Models on Sex
5.4 Sample Weights vs Optimization under Constraint
How can we explain that GridSearch achieves the fairness goal of Demographic Parity, whereas it is not the case of FairDream? The answer involves a more detailed comparison between the models. As introduced above, they differ in the way they implement the fairness constraint:
- In GridSearch Loss, the fairness objective $\mathcal{F}\left(\phi_{\theta,n}\right)$ is a new term added inside the global loss function (this way of correcting is referred to as an optimization of model’s parameters subject to a fairness constraint $\mathcal{F}$ [Wan et al., 2023]). The left part of $L$ is a classic loss function, increasing when the error between the prediction $\widehat{y_{i}}$ and the true $y_{i}$ differs. But the right part is a fairness loss function, added to the standard measurement of error, increasing when the fairness constraint $\mathcal{F}$ is violated. To achieve Demographic Parity, the fairness term of the loss incentivizes the model to predict $\widehat{y_{i}}=1$, even if the true label of a 17-29 year-old individual $i$ is $y_{i}=0$.
- However, FairDream  has no such internal constraint inside its cost function $L(\phi_{\theta,n},\ W_{n})$. To equalize the percentage of individuals predicted as earning over \$50,000, which is the purpose of Demographic Parity, the harmed individuals of the group $S_{k}$ are provided with a new weight of error $w_{n}(S_{k})$, growing with their previous disadvantage. However, once the new sample weights are computed, they become coefficients of the classic loss function in FairDream Loss.
For example, the new weight $w_{2}(S_{1})=0.8$ of any 17-29-year old individual $i$ in the initial observation 3 means that if a 17-29-year old $i$ has a true label of ‘earning less than \$50,000’ ($y_{i}=0$), predicting her as earning over \$50,000 ($\widehat{y_{i}}=1$) amounts to 8 times the cost of any prediction error for 29-37 years old. The FairDream-style classifier is, therefore, strongly incentivized to give predictions in accordance with the true label $y_{i}\in\{0,1\}$. As a consequence, the model becomes more accurate across age groups, by predicting $y_{i}=1$ given their true labels. Which makes the model tend towards Equalized Odds (equality of true and false positive rates).
Although Demographic Parity is set as the fairness purpose, a FairDream-style correction method (here, in-processing reweighting) remains conservative. It does increase the overall positive rates of harmed populations with new weights, but mostly relying on the true labels.
6 Choosing a Fairness Metric conditioned on True Labels
We provided a computational and mathematical explanation for the deviation between the initial fairness objective (Demographic Parity) and the one actually achieved by FairDream (Equalized Odds). In this section we propose a more normative discussion of the choice between these two fairness objectives. As announced in the introduction, Equalized Odds is also commonly opposed to various notions of predictive parity, understood either in terms of equal calibration ([Pleiss et al., 2017, Barocas et al., 2023]), or in terms of equal precision relative to a threshold ([Dieterich et al., 2016, Mayson, 2018]). So first, we start with some data relative to FairDream’s performance on calibration, and we explain why, on a normative basis, calibration for us does not constitute a relevant criterion. The real issue, according to us, is whether the achievement of Equalized Odds remains too conservative compared to the achievement of Demographic Parity.
6.1 Calibration and Precision
One common view of fairness is found in the idea of predictive parity between groups. Predictive parity can mean that the same fraction of true positives is predicted between groups at identical risk scores, or it can mean that the same fraction of true positives is predicted between groups relative to a classification threshold. The first notion corresponds to the statistical notion of calibration of a classifier, whereas the second corresponds to the notion of statistical precision of a classifier. Both notions are sometimes indistinctly referred to under the term “calibration” in the literature (see [Mayson, 2018, Long, 2021]), although strictly speaking they differ. Both notions have also been invoked in defense of COMPAS against ProPublica’s charges of racial bias. In their response to ProPublica, Northpointe argued that at the cutoff chosen by ProPublica to show Unequal Odds between Whites and Blacks, the statistical precision of the COMPAS algorithm was roughly the same between both groups. Moreover, it was shown that at each risk score, the proportion of rearrest between groups was roughly the same (see [Corbett-Davies et al., 2017, Figure 2]),141414[Dieterich et al., 2016] produce very closely related data (see their tables A1 and A2), but they look at predictions above all possible risk scores, which is not strictly speaking calibration. this time evidencing calibration between groups.
Both calibration and precision are central notions in that regard, and one may wonder how FairDream fares with regard to either. Let us consider calibration first. Calibration can be looked at either in a relative, or in an absolute sense ([Eva, 2022]). In a relative sense, it requires that for every possible risk score, the ratios of true positives be equal across groups. That is:
$p(Y=1|R=r,A=1)=p(Y=1|R=r,B=1)$
We select this relative sense as it is the one commonly adopted to link fairness and calibration (see [Corbett-Davies et al., 2017, Hedden, 2021, Barocas et al., 2023]). This sense of calibration has been called “calibration within groups” by [Kleinberg et al., 2016]. Here we simply call it “Equal Calibration”, to indicate that the measure is comparative. A distinct, non-relative sense of calibration, requires that the method used to classify be itself well-calibrated within a given group this time, that is, that for each possible risk score $r$, the ratio of people classified as positive (Y=1) be equal to $r$, or sufficiently close to $r$. This is a stronger requirement, for a method may fail to be well-calibrated in that absolute sense, but still be such that it selects the same percentage of true cases in different groups.
It is a fact that in general, achieving Equalized Odds implies sacrificing Equal Calibration (see [Kleinberg et al., 2016],[Barocas et al., 2023]). The case of the random forest models used above illustrates this trade-off. When we compare and sum the scores (here, probabilities of getting $Y=1$), the calibration gap between groups is increased by FairDream compared to the Baseline model. The area between calibration curves, approximated through a trapezoidal rule, is larger for FairDream than for the initial model (0.2 versus 0.09, Cf Figure 6). That FairDream favors Equalized Odds at the expense of Equal Calibration is also confirmed in the other cases. Out of 16 features where correction happened, FairDream reached the highest gaps between groups in 11 calibration curves (Cf. Table 3.)
Table 3: Comparison between Baseline and FairDream on benchmark: each cell reports the number of times a model achieves the highest gap between groups on a metric
Max Gap between groups Calibration ROC-AUC PR-AUC OPR FPR TPR All metrics
Baseline 5 3 13 9 10 14 54
FairDream 11 11 1 7 6 2 38
Figure 6: Calibration Curves of FairDream and Baseline Random Forest Models on Sex
Figure 6 also tells us something about absolute calibration, so on the positions of curves relative to the ideal calibration curve $x=y$. In the Baseline model, both curves are below ideal calibration: $P(\hat{Y}=1)$ under-estimates the real percentage of women and men which truly earn over \$50,000. Whereas FairDream reverses this tendency: individuals of both groups are over-estimated by the model. However, while switching from under- to over-estimation, FairDream produces a better calibrated curve for women than men, this time judging by how much the female group’s curve departs from absolute calibration. That is, while FairDream undeniably widens the gaps between women and men to fulfill Equalized Odds, the situation of women is enhanced not only on true positive rates, but even on the calibration picture. Overall, therefore, neither the base model nor FairDream is well calibrated in an absolute sense, but they err in opposite ways. Moreover, while the calibration gap is increased with FairDream, this is a case where absolute calibration in the discriminated group is improved in FairDream.
We do not see Equal Calibration to be a fundamental fairness constraint, however, and we do not see it on a par with Equalized Odds either. Indeed, like Demographic Parity, Equalized Odds supposes that a threshold has already been applied to the scores. The decision-maker is only able to act once this threshold has been set. [^15] Like [Grant, 2023], we thus favor decision procedures over predictive methods to investigate the fairness of algorithm. Our defense of Equalized Odds differs from his, but both share this normative consideration.
Yet, as we see in Figure 6, calibration compares the model’s properties along all scores, before any threshold has been set. As argued by [Corbett-Davies et al., 2017], moreover, one can find cases of Equal Calibration between groups for which a decision threshold is nonetheless prone to induce a differential treatment between groups, if the two groups’ risk scores overlap on only part of the scale.
A way to take into account decision thresholds is to compare Equalized Odds with a constraint of Equal Precision between groups, relative to the same threshold, namely:
$p(Y=1|\hat{Y}=1,A=1)=p(Y=1|\hat{Y}=1,B=1)$
Precision, namely the rate of true positives among all predicted positives, is a metric that Northpointe opposed to COMPAS in their rejoinder ([Dieterich et al., 2016]). More generally, [Long, 2021] has argued that Equal Precision ought to stand as a necessary condition on fairness. However, we find it hard to endorse this generalization for all cases. For instance, to use an example of the same kind used by Long, we can come up with cases in which Precision is not equal between two groups, but Sensitivity (= True Positive Rate, aka. Recall) and Selectivity (=True Negative Rate) are identical, and for which we do not have the intuition that there is an unfair treatment.
An example is given in Table 4. Group 1 has 28 students, and Group 2 has 40 students, comprised of students deserving a High Grade and students deserving a Low Grade (where we assume these qualities to be objectively measurable). In both groups, the True Positive Rate and the True Negative Rate are the same: Group 1 and Group 2 have the same absolute number of true High Grade students, but in Group 2 we get more Low grade students wrongly classified as “High” than in Group 1. The Precision on the “High” label is only 1/7 compared to 1/5 in the first group. Since Selectivity and Sensitivity are identical in this example, this is, however, a case in which the criterion of Equalized Odds is satisfied.
Table 4: Unequal Precision but equal Sensitivity and Selectivity ($Y=1$ means that the student’s type is High, $Y=0$ that is is Low)
Group 1 $\hat{Y}=1$ $\hat{Y}=0$
$Y=1$ 1 3
$Y$=0 4 20
Group 2 $\hat{Y}=1$ $\hat{Y}=0$
$Y=1$ 1 3
$Y=0$ 6 30
Against sufficiency this time, we can easily find cases in which Precision on one label will be identical across groups, but for which we would have the intuition of the classifier being unfair or biased. Imagine a teacher having to grade only excellent, High profile students, male and female, in a testing experiment in which the teacher can issue either “High” or “Low” judgments. Suppose that the teacher always grades true High type male students perfectly as “High”, but classifies only 1 in 5 female students as “High” and the others as “Low” (see Table 5). The precision on the “High” predictions is 100 percent in both groups, but obviously the Sensitivity is not the same for males and for females. This is a clear case of a biased judgment, where so many misses in the female group imply an unfair outcome.
Table 5: Equal Precision, Unequal Sensitivity
Males $\hat{Y}=1$ $\hat{Y}=0$
$Y=1$ 25 0
$Y$=0 0 0
Females $\hat{Y}=1$ $\hat{Y}=0$
$Y=1$ 5 20
$Y=0$ 0 0
The upshot is that Equal Precision is neither a necessary, nor a sufficient condition on fairness. In the case we just presented, one might respond that Precision in the “Low” judgments is not constant across the groups, even as precision on the “High” judgments is. But the problem is that to ask for Equal Precision on both labels is tantamount to asking for Equalized Odds, which would undercut Long’s argument. This is not to say that Precision does not matter to FairDream. As explained in the previous section (see fn. 12), we conducted our benchmark so as to maximize the trade-off between Precision and Recall. From Table 3, we also see that the gap in PR-AUC diminishes in FairDream across groups compared to the baseline. But since this is not a criterion we set a priori, no more than ROC-AUC, we do not build any conclusion on this fact.
In summary, therefore, whether on a descriptive basis, or on a normative basis, Equal Calibration does not end up as a decisive criterion in our evaluation of FairDream’s correction, and neither does Equal Precision. In what follows, therefore, we return to the opposition between Equalized Odds and Demographic Parity as opposite requirements.
6.2 Fairness from Accuracy
In our benchmark comparison, each time the GridSearch models outperformed FairDream with regard to Demographic Parity, it was at the cost of a worse statistical performance. [^16] Cf. Appendix B. On the contrary, FairDream gives a truthful picture of the ground labels: the FairDream correction method performs better in terms of accuracy, with a low rate of false positives (less than 5%) and a high rate of true positives (nearly 90%).
The question raised by this situation is whether increasing statistical accuracy, namely Selectivity and Sensitivity, is automatically a way of increasing fairness. Several arguments can be given against the sufficiency of accuracy to achieve fairness. The main one is that even the most accurate classification may simply replicate imbalances that are in the data ahead of the algorithm’s workings, as a result of social biases or social injustice.
We agree with this argument, and we grant that improving on descriptive accuracy may not be sufficient to achieve normative fairness (in line with Hume’s classic remarks on the is-ought distinction [Hume, 1739]). However, we consider that descriptive accuracy in predictions should at least be viewed as a necessary condition on fairness [Wachter et al., 2020], and that it may even come out necessary and sufficient in cases in which the predictions themselves concern matters of fact.
In matters of fact, the case for conditional fairness metrics can be buttressed by general epistemological considerations. A user trying to narrow the gap between overall positive rates made on age groups by our FairDream classifier (16% for 17-29 years old and 59% for 29-37 years old, Cf. Table 2) may be viewed as falling prey to a “base rate neglect” ([Tversky et al., 1982]). It is a common mistake when interpreting statistics to overlook the base rates. In the present case, the base rate, or the ratio of actual positives to the whole population, is radically different in each age group. Less than one in six younger individuals actually earns over \$50,000, while four in six older actually earn over \$50,000.171717The base rates are 547/3616=15% for 17-29 years population, versus 2400/3666=65% for 29-37 years population (Cf. Table 2).
Provided that the data is reliable, it is therefore not possible for a classifier to predict more than 15% of the 17-29 years old and 65% of the 29-37 years old individuals as earning over \$50,000 without increasing false positives (thereby decreasing in predictive performance). The fact that the model maximized the true positive rates and minimized the false positive rates pinpoints that the accuracy by group is equalized by FairDream. The individuals predicted to be positives are now closer to their base rates. [^18] We do not claim here that the equality of base rates is a condition for algorithmic fairness. As [Hedden, 2021] shows, base rates can be equal while the classifier is unfair. We only point that in cases where true labels are agreed upon, when the prediction of positives is close to base rates among previously harmed and not harmed groups, the equality of true and false positive rates indicates a maximally possible fair situation.
Further arguments in favor of accuracy concern downstream effects (see [Barocas et al., 2023]): blindly enforcing Demographic Parity by increasing the false positive rate in the younger group is also susceptible to lead to harmful consequences. Ceteris paribus, it can put the younger group at higher risk of defaulting their loan, assuming income is causally the main variable to sustain loan refunding, and that the algorithm’s prediction on income is critical in deciding whether or not to grant credit.
That being said, the method’s emphasis on Equalized Odds is not as conservative as it might seem. In particular, it may be objected that if Sensitivity and Specificity are key here, then they should be plotted for each decision threshold, which is what ROC-AUC represents. But as Table 3 indicates, in the benchmark study the gap between ROC-AUC is increased on average compared to the baseline (whereas for PR-AUC it is decreased).
The situation is reminiscent here of the debate between ProPublica and Northpointe: both of them, after all, agree that both Selectivity and Sensitivity matter for fairness, but Northpointe’s claim is that the tradeoff between them should be measured at all thresholds by comparing the ROC AUC, and not relative to just one threshold. However, Northpointe also grants that given a particular decision boundary, one may compare true positive rates and false positive rates between groups to check for inequalities in accuracy. Their proposed criterion is that if the ratios between TPR is less than 1 at some threshold, and the ratio of FPR is greater than 1, then one can conclude for inequality, though they concede that in cases in which only one condition is satisfied, then cost considerations may be brought in. [^19] See [Dieterich et al., 2016, p. 16]: “The [Ratio of FPR] is less than 1, but the [Ratio of TPR] is not greater than 1. This is an inconclusive result. When the results of a comparison of the accuracy of a binary test in two groups is inconclusive, the results can be consolidated using a cost function (Pepe, 2003). That work is beyond the scope our report”. The situation depicted in Table 1 has this property: there, the imbalance exceeds 1 only in the false positive rate. However, it does it by a factor 10 (0.2 in the 29-37 yr-old group compared to 0.02 in the 17-29 yr-old group). Even for those who sustain that COMPAS is not biased, this is a case were the imbalance justifies intervention.
As shown in Table 2, the proposed intervention by FairDream does narrow the gap in Demographic parity. The fact that it does this cautiously, namely within the bounds of Equalized Odds, still corresponds to a policy that is less conservative than one that would suggest a revision only in cases in which the ratios of TPR and FPR both exceed the threshold of 1 discussed by Northpointe in particular.
6.3 Always prefer Equalized Odds?
The foregoing arguments notwithstanding, there remain reasons to doubt the validity of conditional metrics in every situation. In order to answer this worry, we use the same methodology, namely we need to vary the case and to consider more examples. For comparison, consider another realistic financial classification task in which the bank-teller has to estimate which client is a “good risk” or “bad risk” based on age, sex, profession, accounts, and credit amount, as appears in the German Credit dataset. [^20] https://www.kaggle.com/datasets/uciml/german-credit.
Here, a hidden bias may very well have influenced the decision to label clients as “good risk” or “bad risk”, simply because the labels “good” and “bad” are not purely factual (unlike “earning over \$50,000”), but evaluative and judge-dependent ([Solt, 2018]). The bias may be reflected in spurious correlations between variables, for instance between “good risk” and male. Applied to this distribution of the labels, a model achieving Equalized Odds will predict overall positive rates which replicate wrong or undesirable base rates, under-representing the “good risk” women.
The choice between conditional and unconditional fairness criteria represents two different visions of distributive justice ([Rawls, 1999]). An appeal to unconditional fairness criteria is justified when the basic labels and their distribution can be suspected of an initial unwanted bias [Wachter et al., 2020]. On the other hand, the use of conditional metrics is justified when the goal is to ground prediction in descriptions and distributional properties of a population that need to be faithfully represented [Dwork et al., 2012].
In our running example of income prediction, the transparency of labels and their factual character is undeniable. The distribution of income levels, on the other hand, may be an object of deeper social debate (viz. should the youth’s first income be raised?). [^21] See for example [Taylor, 2020] But merely changing the labels would not appear to be an adequate or efficient way of affecting the underlying distribution of incomes. In a task involving the labels of “good” vs “bad” risk clients, on the other hand, those very labels are no longer transparent, and the various methods of correction we distinguished (pre-processing, in-processing, and post-processing) may be used to remove hidden biases.
Acting on the labels is a delicate matter, which can only be justified if the labels are evaluative, even implicitly, and if there are associated leverages to enforce the new distribution of resources (public policies, internal rules and means inside corporations). In particular, it seems to us that an understanding of causal relations is required to identify the features on which to correct the labels, in order not to harm new individuals. For example, the investigation could focus on the effects on changing the risk labels on the basis of a feature like income, or geographic area. [^22] As a first step, the effects of overall fairness metrics on new allocation of outcomes could be validated with A/B testing on real populations ([Saint-Jacques et al., 2020]). Controlled experimentation on the effects of new labels distribution would defuse two problems of algorithmic fairness spotted by [Selbst et al., 2019]. Providing contextualization on the impact of new labels against the “portability trap”, it would avoid such a reallocation to be counterproductive as the “ripple effect trap” pinpoints.
It is important to bear in mind that FairDream was implemented to inform users of disparities in overall positive predictions between various groups. It is possible, of course, to imagine cases where a gap in overall positive predictions is detected even as the True positive rate and the False positive rate are equal between two groups from the get go, and in cases for which sensitivity and selectivity are actually perfect. For such cases, which imply different base rates between the groups, we are not saying that no intervention should ever be made. The Census case we started off with is of this kind: on a variant of it, each adult in their group is perfectly predicted for their income, and the difference in overall positives is merely due to the fact that, on average, younger adults earn less than the older cohorts. FairDream in such cases will not offer to close this gap, but it will still be instrumental in providing information [Mayson, 2018] about inequalities that may be structural.
7 Conclusion
In this paper we implemented an algorithmic fairness tool to detect and correct for disparities in classification, the package FairDream. The practical goal of this package ‘is to help lay users implement their vision of fairness, in order to come up with socially optimal decisions. Our study of the specific properties of FairDream raised a puzzle and an explainability issue: how can an algorithm designed to enforce Demographic Parity end up converging on a different fairness goal, namely Equalized Odds?
In the first part of this paper, we provided an in-depth explanation of this specific property of FairDream. We conducted a systematic benchmark to compare FairDream with the state-of-the-art correction method GridSearch, testing the two methods on different features and machine-learning models. While the two methods are in-processing correction approaches, we found that GridSearch is suitable for achieving Demographic Parity, whereas FairDream’s correction method results in Equalized Odds. Still, as the example given in Table 2 shows us, the results produced by FairDream do narrow the gap in Demographic Parity between the groups of interest, though within the bounds of reaching Equalized Odds.
First, we proposed an explanation for this difference by looking at the specifics of each algorithm. GridSearch  incorporates the fairness constraint as an additional cost inside the cost function of the model. Due to the fairness cost, this family of methods is more likely to deviate from the true labels (increasing false positives), in order to achieve Demographic Parity. On the other hand, FairDream introduces a differentiated attention to harmed individuals inside the cost function of the model. Therefore, it increases accuracy on harmed groups (through enhanced true positives). By basing fairness on accuracy, FairDream-type methods are thus bound to make the overall positive rates converge towards their base rates. As a result, they will rarely achieve Demographic Parity. But the method, we have also argued, also leads us to the conclusion that Equalized Odds is more relevant than either Equal calibration or Equal Precision in relation to fairness intervention.
Our study of FairDream also led us to a normative conclusion about how to approach fairness. By choosing a conditional metric, the user implicitly commits to the idea that there is sufficient agreement on the transparency of labels and on the shape of their distribution; by selecting an unconditional metric, one commits to the idea that the labels themselves may have introduced unwanted disadvantages, or that the input distribution must be corrected rather than replicated. In the case of FairDream, the default is therefore set to a correction method that may be viewed as overly prudent, but this is as it should be, since as we argued, more revisionary methods need to be carefully justified and cannot be applied lightly.
Finally, our investigation of FairDream allows us to draw a more general lesson concerning the explainability of AI models. Prima facie, FairDream may be seen as operationalizing a simple interventionist fairness metric, namely Demographic Parity. In practice, however, it produces results that are significantly different and that enforce Equalized Odds. As we have argued, the corrections produced by FairDream can be rationalized and normatively justified, but this involves an explanation of the reweighting method used by the algorithm, and a contrastive look at nearby systems governed by similar correction mechanisms that produce different outcomes. We conclude that AI classification systems should preferably be released with an indication of these features: which goal they seek to operationalize, how they actually perform relative to this goal, and finally how they perform on alternative data, and relative to minimally modified variants on the same data. Making this information available dispels arbitrariness in classification, and is a step toward more trust and more control on the side of the users.
Appendix
A FairDream’s Reweighting
We provide here more details on how the sample weights are generated during the in-processing correction of FairDream. The differentiated weights of error are established in a deterministic but simple way, to grant user’s understanding on the process of correction for fairness. For a FairDream model $n$in $[1,5]$ of five models to be trained in competition, the new weight of error for each group ${S_{k}}$ is computed that way:
$w_{n}(S_{k})={\tt rate_{\_}indivs_{\_}disadvantaged}(S_{k})\times exp(n\times{\tt gap_{\_}fair_{\_}scores}(S_{k}))$
The exponential part of the equation focuses on how far from the maximum the current fairness score (e.g. overall positive rate) of the group $S_{k}$ is. It resets the weight according to the disadvantage of $S_{k}$:
${\tt gap\_fair\_scores}(S_{k})=|{\tt fair\_score}(S_{k})-max({\tt fair\_score}(S_{1}),..,{\tt fair\_score}(S_{n}))|$
The new weight $w_{n}(S_{k})$ takes into account the difference across fairness scores, but also the number of individuals previously impacted. ${\tt rate\_indivs\_disadvantaged}$ is a coefficient which stresses the number of people disadvantaged inside the group, relative to the overall population. This coefficient increases as the people in $S_{k}$ are a higher share of the population:
${\tt rate\_indivs\_disadvantaged}(S_{k})={\tt gap\_fair\_scores}(S_{k})\times\frac{|S_{k}|}{\sum_{i=1}^{n}|{S_{i}}|}$
B Benchmark on Census - models and features
Table 6: Results of the benchmark of GridSearch and FairDream to equalize overall positive rates (OPR), with true positive rates (TPR) false positive rates (FPR), AUCs (ROC and PR), and calibration errors (as area to perfectly calibrated curve)
Max Gap between groups OPR FPR TPR ROC-AUC PR-AUC Calibration All
GridSearch 7 9 10 11 12 8 57
FairDream 9 7 6 3 2 8 35
In this appendix, we synthesize the results of the experiment described in Section 5.3. Table 6 displays, aggregated for all models and all features where a correction process was initiated  [^23] We do not count here the results on neural networks which were non significant, as FairDream and GridSearch systematically implemented random classifiers to reach perfect Demographic Parity. See the repository for more details. the performances of GridSearch and FairDream when set the task of equalizing overall positive rates (corresponding to Demographic Parity).
Each model received one point (of penalty) when it achieved the highest gap between the groups with the minimal and maximal scores. In Table 6, lower scores indicate better results on that count. To verify the convergence property of FairDream to Equalized Odds, the table also indicates the number of times a model was worse than the other at equalizing true positive rates, and false positive rates, across the relevant features.
The table shows that GridSearch slightly outperformed FairDream to achieve Demographic Parity. In total, we find GridSearch to better fulfill Demographic Parity than FairDream on 9/16 cases (where FairDream meets the highest gaps between groups).
However, this is at the cost of a decrease in predictive performance. On both AUCs, GridSearch digs the worst gaps between groups in almost every case. Besides, the mean ROC-AUC of all GridSearch models is 59%, which is a loss of 20% compared to the baseline classifier (79%). The low ROC-AUC of GridSearch, as well as the plots of the repository, indicate that Demographic Parity is achieved by increasing true positive rates, but thereby wrongly enhancing the false positive rates.
While GridSearch outperforms FairDream on equalizing overall positive rates, FairDream achieves better results to equalize false positive rates (9/16 cases) and true positive rates (10/16 cases), with a ROC-AUC that remains high ($78\%\mp 1\%$ compared with the baseline model). This highlights that even if set to respect Demographic Parity, FairDream increases true positive rates and decreases false positive rates, confirmed by the plots of the repository https://anonymous.4open.science/r/weights_distortion_impact-15/.